OpenStudy (anonymous):

Can someone help me with this please? 6+sqr24/2. the answer is 3+sqr6 but I am getting 3sqr12. *sqr = radical sign

OpenStudy (anonymous):

What you have is [6 + sqr(24)]/2. What you evaluated is 6 + sqr(24/2). Do you see the difference? Parentheses are important. You have to be careful. First you need to simplify sqr(24) into 2sqr(6). Then take your numerator 6 + 2 sqr(6) and factor out a 2: 2[3 + sqr(6)]. Now you can cancel the 2 you just factored out with the one in the denominator and get the answer 3 + sqr(6).

OpenStudy (anonymous):

I don't understand why the sqr of 24 is implified to 2 times the sqr of 6.

OpenStudy (anonymous):

simplified*

OpenStudy (anonymous):

24 = 4 * 6 24 = 2^2 * 6 [because 4 = 2^2] sqrt(24) = sqrt(2^2 * 6) sqrt(24) = sqrt(2^2) * sqrt(6) [because sqrt(AB) = sqrt(A)*sqrt(B) sqrt(24) = 2 * sqrt(6) Is that what you needed? The reason you need to simplify it first is that in order to simplify the entire expression, you will need to cancel the 2 in the denominator with a factor of 2 in the numerator. To do that, you have to have the entire numerator factored, not just part of it. For instance you can't simplify (6x + 8)/2 to 3x + 8 or 6x + 4. You have to factor the 6x + 8 first. (6x + 8)/2 = 2(3x + 4)/2 = 3x + 4. Your problem is similar, but it is a bit more complicated because of the square root in it.

OpenStudy (anonymous):

okay, thank you :)